### v1

# A Scalable Measure of Loss Landscape Curvature for Analyzing the Training Dynamics of LLMs

Abstract
Understanding the curvature evolution of the loss landscape is fundamental to analyzing the training dynamics of neural networks. The most commonly studied measure, Hessian sharpness ($\lambda_{\max}^{H}$) —the largest eigenvalue of the loss Hessian —determines local training stability and interacts with the learning rate throughout training. Despite its significance in analyzing training dynamics, direct measurement of Hessian sharpness remains prohibitive for Large Language Models (LLMs) due to high computational cost. We analyze critical sharpness ($\lambda_{c}$), a computationally efficient measure requiring fewer than $10$ forward passes given the update direction $\Delta\bm{\theta}$. Critically, this measure captures well-documented Hessian sharpness phenomena, including progressive sharpening and Edge of Stability. Using this measure, we provide the first demonstration of these sharpness phenomena at scale, up to 77B parameters, spanning both pre-training and mid-training of OLMo-2 models. We further introduce relative critical sharpness ($\lambda_{c}^{1\to 2}$), which quantifies the curvature of one loss landscape while optimizing another, to analyze the transition from pre-training to fine-tuning and guide data mixing strategies. Critical sharpness provides practitioners with a practical tool for diagnosing curvature dynamics and informing data composition choices at scale. More broadly, our work shows that scalable curvature measures can provide actionable insights for large-scale training.
[…]
1 Introduction
Understanding the evolution of the loss landscape $L(\bm{\theta})$ over the high-dimensional parameter space $\bm{\theta}$ is fundamental to analyzing the training dynamics of neural networks. The loss landscape describes how the objective function changes with model parameters, and its geometry directly influences optimization, generalization and stability throughout training (Gilmer et al., 2022; Kalra and Barkeshli, 2024). Intuitively, gradient-based optimization methods can swiftly navigate to the minima in smooth landscapes, while for rough landscapes, such methods can get trapped, hindering convergence or potentially converging to suboptimal solutions (Li et al., 2018).
The local geometry of the loss landscape is commonly examined through the eigenvalues and eigenvectors of its Hessian matrix $H(\bm{\theta}):=\nabla_{\theta}^{2}L(\bm{\theta})$ (Wu et al., 2018; Lewkowycz et al., 2020; Cohen et al., 2021). These eigenvalues provide insights into the local curvature of the loss landscape —large eigenvalues correspond to sharp, steep directions, while small eigenvalues indicate flat, smooth regions. The largest eigenvalue of the Hessian, which is often referred to as Hessian sharpness $\lambda^{H}_{\max}$, quantifies the worst-case curvature of the landscape and is a fundamental metric in optimization. Its reciprocal, flatness $1/\lambda_{\max}^{H}$, is a complementary measure used to describe the curvature. In neural network optimization, Hessian sharpness relates to the local training stability (Wu et al., 2018; Lewkowycz et al., 2020). For vanilla Gradient Descent (GD), if the learning rate exceeds the threshold $\sim 2/{\lambda_{\max}^{H}}$, loss increases as training ‘catapults’ out of the local basin and eventually converges to a flatter region where the stability conditions are satisfied (Lewkowycz et al., 2020). This observation generalizes to more complex optimization problems, including mini-batch settings and adaptive optimizers, albeit governed by different notions of sharpness (Wu et al., 2018; Agarwala and Pennington, 2025; Cohen et al., 2024; Kalra and Barkeshli, 2024).
Figure 1: Hessian sharpness exhibits progressive sharpening and Edge of Stability (EoS) under constant learning rate. The dashed lines corresponding to the learning rate mark the EoS threshold.
As illustrated in Figure˜1, Hessian sharpness exhibits several robust trends throughout neural network training, particularly with constant learning rates. Training typically begins with an early reduction in Hessian sharpness (Kalra and Barkeshli, 2023; Kalra et al., 2025) followed by a continuous increase until it reaches the stability threshold (Jastrzebski et al., 2020; Cohen et al., 2021). Once this threshold is reached, the training stabilizes through a self-stabilization mechanism (Damian et al., 2023; Cohen et al., 2025) —instead of diverging as classical optimization would predict, Hessian sharpness begins to oscillate around this critical value. The continual increase in Hessian sharpness is referred to as progressive sharpening while the subsequent oscillations around the critical threshold is termed the Edge of Stability (EoS) (Cohen et al., 2021). For the more realistic setting of learning rate schedules involving warmup and decay, sharpness closely follows the learning rate schedule (Gilmer et al., 2022; Kalra and Barkeshli, 2023; Cohen et al., 2021). Due to its particularly close relationship with learning rate, sharpness can be used as a diagnostic tool for identifying training instabilities.
Beyond training dynamics, sharpness of the final solutions is linked to generalization properties, motivated by the intuition that flatter minima generalize better (Hochreiter and Schmidhuber, 1997). However, this relationship has been called into question (Kaur et al., 2023), as empirical analyses show mixed results —while flatter solutions generalize better in some settings, they can hurt performance in others. Sharpness also influences the early phase of training, with flatter initializations typically achieving better performance (Dauphin and Schoenholz, 2019; Kalra and Barkeshli, 2024).
Despite its value for examining model initialization, training dynamics, and generalization properties, analyzing sharpness at scale is challenging. Computing Hessian sharpness relies on iterative eigenvalue solvers (e.g., Power iteration, Lanczos), which require repeated Hessian-vector products (HVPs). However, HVP computation via automatic differentiation is often incompatible with modern training efficiency tools; kernels like Flash Attention (Dao et al., 2022) typically lack second-derivative implementations required for double backpropagation. Furthermore, iterative solvers can require hundreds of iterations with potential convergence failures. As a result, most existing studies are restricted to small-scale experiments (typically $\sim 10$M parameters), leaving open questions about how sharpness evolves in Large Language Models (LLMs) at scale, and how it relates to optimization and downstream performance. In this work, we address this challenge by analyzing critical sharpness, which leverages the relationship between curvature and training instability to serve as a computationally efficient proxy for the loss landscape curvature of LLMs.
1.1 Contributions
Our contributions are as follows:
  • We analyze critical sharpness, defined as $\lambda_{c}=2/\eta_{c}$, where $\eta_{c}$ is the critical learning rate—the smallest learning rate that causes the training loss to increase in the next training step. To estimate the critical learning rate $\eta_{c}$, we perform an efficient line search along the update direction $\Delta\bm{\theta}$ from training. This procedure only requires forward passes, making it fully compatible with modern large-scale distributed training infrastructure, while avoiding the convergence issues of iterative eigenvalue solvers. In practice, we find that this procedure reliably estimates critical sharpness in only 55-66 forward passes111except for the first iteration, which depends on the initial guess $\eta_{0}$ for line search..
  • We then examine the relationship between critical sharpness and Hessian sharpness. Under the quadratic loss approximation, we show that critical sharpness can be written as a weighted sum of Hessian eigenvalues and the two measures coincide when the gradient aligns with the largest eigenvector of the Hessian. We also generalize this result to adaptive optimizers.
  • We demonstrate that critical sharpness reliably captures well-documented Hessian sharpness phenomena, such as progressive sharpening and the Edge of Stability (Cohen et al., 2021). This makes critical sharpness a practical and computationally efficient proxy for sharpness. Leveraging OLMo-2 checkpoints (Walsh et al., 2025) across pre-training and mid-training, we demonstrate progressive sharpening persists at realistic scales throughout training, including models with up to 7 billion parameters.
  • We introduce a relative measure of critical sharpness $\lambda_{c}^{1\to 2}$ to quantify the curvature of one loss relative to another, with application to analyzing the pre-training loss landscape during mid-training. By varying the mix ratio of pre-training and fine-tuning (e.g. math), we show that increasing the proportion of pre-training data lowers relative critical sharpness, helping the model stay near the “pre-trained basin” during finetuning. We further demonstrate that downstream performance depends on basin retention: GSM8K (Cobbe et al., 2021) (math) benefits from leaving the pre-trained basin ($\eta>2/\lambda_{c}^{1\to 2}$), while MMLU (Hendrycks et al., 2021) (generic reasoning) performs better when the basin is retained ($\eta<2/\lambda_{c}^{1\to 2}$). This enables us to prescribe data mixing strategies to balance performance on math and generic reasoning tasks, depending on the desired outcome.
2 Characterizing Critical Sharpness: Theory and Empirics
In this section, we characterize the dynamics of critical sharpness both theoretically and empirically. We first establish the theoretical relationship between critical sharpness and Hessian sharpness under a quadratic loss approximation, then validate these insights empirically on FCNs trained on CIFAR-10 using SGD across different batch sizes (Figure˜3).
2.1 Setup
Consider a model with trainable parameters $\bm{\theta}\in\mathbb{R}^{n}$, and let $L(\bm{\theta})$ denote the loss function with gradient $g(\bm{\theta}):=\nabla_{\theta}L(\bm{\theta})$ and Hessian $H(\bm{\theta}):=\nabla_{\theta}^{2}L(\bm{\theta})$. Let $\{\lambda_{i}^{H}\}_{i=1}^{n}$ and $\{{u}_{i}\}_{i=1}^{n}$ denote the eigenvalues and eigenvectors of the Hessian, respectively. The largest eigenvalue $\lambda_{\max}^{H}=\max_{i}\lambda_{i}^{H}$, termed Hessian sharpness, quantifies the worst-case local curvature. For adaptive optimizers with pre-conditioner $P(\bm{\theta})$ (e.g., Adam), the training dynamics is governed by the pre-conditioned Hessian $P(\bm{\theta})^{-1/2}H(\bm{\theta})P(\bm{\theta})^{-1/2}$, with eigenvalues $\{\lambda^{PH}_{i}\}_{i=1}^{n}$ and eigenvectors $\{{v}_{i}\}_{i=1}^{n}$. We omit explicit $\bm{\theta}$ dependence when clear from context.
Figure 2: Comparison of different sharpness measures on an illustrative landscape featuring a sharp valley direction and a flat river direction (Wen et al., 2025). Hessian sharpness quantifies the curvature along the sharpest direction of the landscape (the valley). Directional sharpness measures the quadratic curvature along $\Delta\bm{\theta}$, while critical sharpness, its empirical counterpart, quantifies how far one can step along $\Delta\bm{\theta}$ before the loss increases.
2.2 Critical Sharpness: A Scalable Measure of Curvature
Definition 2.1 (Critical learning rate $\eta_{c}$ and Sharpness $\lambda_{c}$).
Given an update direction $\Delta\bm{\theta}$, we define critical learning rate as the smallest learning rate that causes the loss to increase in the next training step:
$\eta_{c}=\min_{\eta>0}\{\eta\mid L(\bm{\theta}-\eta\Delta\bm{\theta})>L(\bm{\theta})\}.$ (1)
Correspondingly, we define critical sharpness as the scaled reciprocal of the critical learning rate $\lambda_{c}=2/\eta_{c}$.
Geometric Interpretation:
Critical sharpness provides an intuitive, optimizer-aware measure of the local curvature of the loss landscape. Geometrically, it quantifies the “natural length-scale” of the landscape by estimating how far one can move in the current update direction without leaving the local region. While Hessian sharpness quantifies the curvature along the steepest direction of the landscape, critical sharpness measures the curvature along the update direction $\Delta\bm{\theta}$. Figure˜2 compares critical sharpness with Hessian sharpness on an illustrative landscape.
Efficient estimation of critical learning rate:
To efficiently measure the critical learning rate, we build upon the line search method proposed by Kalra and Barkeshli (2024). Given the update direction $\Delta\bm{\theta}$ from training, we compute the critical learning rate $\eta_{c}$ using a two-phase line search procedure:
  1. 1.
    Exponential Search: Starting from an initial guess $\eta_{0}$, we iteratively double (halve) $\eta$ until the loss $L(\bm{\theta}-\eta\Delta\bm{\theta})$ exceeds (or falls below) the current loss $L(\bm{\theta})$. This quickly identifies an interval $\eta_{\text{lower}},\eta_{\text{upper}}]$ containing $\eta_{c}$.
  2. 2.
    Binary Search: We then refine this interval using a binary search until the relative error $|1-\eta_{\text{lower}}/\eta_{\text{upper}}|$ falls below a threshold $\epsilon$. For $\epsilon=\frac{1}{2^{k}}$, the binary search converges in $k$ steps.
We approximate the critical learning rate using the mean of the final range $\eta_{c}\approx\frac{1}{2}(\eta_{\text{lower}}+\eta_{\text{upper}})$. In practice, the exponential search typically requires only 1122 iterations (except for the first use, which depends on the initial guess $\eta_{0}$), and setting $\epsilon=1/16$ (i.e., $k=4$) provides a reliable and efficient estimate of the critical learning rate in approximately 5566 forward passes, as shown in Figure˜3. Our approach is scalable, as it relies solely on forward passes to evaluate $L(\bm{\theta}-\eta\Delta\bm{\theta})$ and leverages the same computational primitives as standard first-order distributed training, avoiding the challenges of Hessian-based methods. In Section˜7, we provide the detailed algorithm and further discuss the design choices.
2.3 The Relationship between Critical Sharpness and Hessian Sharpness
(a) Batch size = 50,000 (GD)
(b) Batch size = 5000
(c) Batch size = 500
Figure 3: Comparison of different sharpness measures for MLPs trained on CIFAR-10 image classification task using SGD with learning rate $\eta$. Both critical and directional sharpness exhibit progressive sharpening and Edge of Stability, albeit some deviations from Hessian sharpness. The dashed line denotes the Edge of Stability threshold, given by $2/\eta$.
Before investigating the empirical behavior of critical sharpness, we establish its theoretical relationship with Hessian sharpness. This connection provides a principled foundation for understanding what critical sharpness captures. To this end, we consider the quadratic approximation of the loss function along $\Delta\bm{\theta}$:
$L(\bm{\theta}-\eta\Delta\bm{\theta})\approx L(\bm{\theta})-\eta\Delta\bm{\theta}^{T}g(\bm{\theta})+\frac{1}{2}\eta^{2}\Delta\bm{\theta}^{T}H(\bm{\theta})\Delta\bm{\theta}.$
Within this approximation, the loss will increase if the learning rate $\eta$ exceeds $2/\lambda_{\text{dir}}$, where $\lambda_{\text{dir}}$ is the directional sharpness (Pan and Li, 2022; Roulet et al., 2024):
Definition 2.2 (Directional Sharpness $\lambda_{\text{dir}}$).
The directional sharpness along the update direction $\Delta\bm{\theta}$ is:
$\lambda_{\text{dir}}=\frac{\Delta\bm{\theta}^{T}H(\bm{\theta})\Delta\bm{\theta}}{\Delta\bm{\theta}^{T}g(\bm{\theta})}.$ (2)
Directional sharpness $\lambda_{\text{dir}}$ serves as an analytically tractable approximation to the empirically measured critical sharpness. Figure˜3 shows that directional sharpness closely tracks critical sharpness throughout training. This alignment suggests that the local quadratic approximation well captures the complex dynamics in this simplistic setting. For Gradient Descent (GD), the directional sharpness can be expressed as a weighted sum of Hessian eigenvalues (Roulet et al., 2024):
Result 2.1 (Relationship between Directional and Hessian Sharpness for Gradient Descent).
For Gradient Descent (GD), the directional sharpness $\lambda_{\text{dir}}$ can be expressed as a weighted sum of the Hessian eigenvalues $\{\lambda_{i}^{H}\}_{i=1}^{n}$, where the weights quantify the alignment of the gradient with Hessian eigendirections $\{u_{i}\}_{i=1}^{n}$:
$\lambda_{\text{dir}}=\frac{\sum_{i=1}^{n}c_{i}^{2}\lambda_{i}^{H}}{\sum_{i=1}^{n}c_{i}^{2}}\leq\lambda_{\max}^{H},$ (3)
where $c_{i}=\bm{g}^{T}\bm{u}_{i}$ is the projection of the gradient onto the $i^{th}$ eigenvector.
It follows that if the gradient $\bm{g}$ is perfectly aligned with the largest eigenvector, i.e., $\bm{g}\propto\bm{u}_{\max}$, then directional sharpness coincides with Hessian sharpness. More generally, directional sharpness is always bounded above by Hessian sharpness. As a consequence, the gap between Hessian and directional sharpness quantifies the alignment of the gradient with the top eigendirection of the Hessian. We now generalize this result to adaptive optimizers.
Result 2.2 (Relationship between Directional and Hessian Sharpness for Adaptive optimizers).
For adaptive optimizers (e.g., Adam) with pre-conditioner $P$, the directional sharpness $\lambda_{\text{dir}}$ can be expressed as a weighted sum of the pre-conditioned Hessian eigenvalues $\{\lambda^{PH}_{i}\}_{i=1}^{n}$, where the weights quantify the alignment of the pre-conditioned gradient $P^{-1/2}\bm{g}$ with pre-conditioned Hessian eigendirections $\{\bm{v}_{i}\}_{i=1}^{n}$:
$\lambda_{\text{dir}}=\frac{\sum_{i=1}^{n}c_{i}^{2}\lambda_{i}^{PH}}{\sum_{i=1}^{n}c_{i}^{2}}\leq\lambda_{\max}^{PH},$ (4)
where $c_{i}=P^{-1/2}\bm{g}^{T}\bm{v}_{i}$ is the projection of the pre-conditioned gradient $P^{-1/2}g(\bm{\theta})$ onto the $i^{th}$ eigenvector $\bm{v}_{i}$ of the pre-conditioned Hessian.
Together, these results establish a connection between directional sharpness and Hessian sharpness, and indicate that the two measures diverge when the alignment between the gradient and the top eigendirection is small. In turn, Hessian and critical sharpness coincide when the loss surface is approximately locally quadratic, and the gradient is primarily along the top eigendirection. We detail the derivations of the above results in Section˜10.1.
We examine the differences between the three sharpness measures in Figure˜3. In the full-batch regime, Hessian sharpness exhibits progressive sharpening, eventually reaching the Edge of Stability (EoS) threshold. In contrast, both directional sharpness and critical sharpness remain nearly constant during the early stages of training, followed by an abrupt increase to the EoS threshold. At smaller batch sizes, however, all three sharpness measures display a gradual increase from the onset of training. Notably, after crossing the EoS threshold, Hessian sharpness tends to oscillate above the threshold, whereas both critical sharpness and directional sharpness oscillate more closely around it. Overall, critical sharpness exhibits progressive sharpening and Edge of Stability, albeit with some differences compared to Hessian sharpness discussed above. In the following section, we extend our analysis to more realistic, large-scale training settings.
3 Critical Sharpness Dynamics at Scale
Modern large-scale models are typically trained using Adam with weight decay (Loshchilov and Hutter, 2019), which helps mitigate training instabilities (D’Angelo et al., 2024). To analyze sharpness dynamics in this context, we first analyze the stability threshold for common optimizers with weight decay:
Result 3.1 (Stability threshold for optimizers with weight decay).
For gradient descent and adaptive optimizers such as Adam, adding weight decay shifts the EoS threshold by a constant that depends on the decay strength $\gamma$:
$\lambda^{H}_{\max}$ $=\frac{2}{\eta}-\gamma\quad$ (GD)
$\lambda^{PH}_{\max}$ $=\left(\frac{2}{\eta}-\gamma\right)\left(\frac{1+\beta_{1}}{1-\beta_{1}}\right)\quad$ (Adam) (5)
where $\eta$ is the learning rate, $\beta_{1}$ is Adam’s momentum parameter, and $\gamma$ is the weight decay strength. We provide the detailed proofs in Section˜10.2.
We now analyze GPT-style Transformers pre-trained for next-token prediction on the FineWebEdu dataset (Penedo et al., 2024) using AdamW with Warmup-Stable-Decay (WSD) schedule (Hu et al., 2024). Figure˜4 compares the dynamics of critical and pre-conditioned sharpness for three different learning rates. The pre-conditioned sharpness $\lambda^{PH}_{\max}$ exhibits progressive sharpening while following the learning rate schedule throughout training —it is pushed down during the warmup phase, stays constant at the stability threshold during the stable phase, and increases again when the learning rate is decayed. Critical sharpness and directional sharpness follow similar trends throughout training —they exhibit clear progressive sharpening and EoS behavior, while oscillating below the EoS threshold during the stable phase and capturing the increase in pre-conditioned sharpness during the decay phase.
Figure 4: Dynamics of pre-conditioned, directional, and critical sharpness during GPT-style Transformer training on FineWebEdu with AdamW using a Warmup-Stable-Decay (WSD) schedule. Critical sharpness tracks pre-conditioned sharpness throughout training, making it an effective proxy. The colored dashed lines denote the theoretical learning rate threshold $(2+2\beta_{1})/(1-\beta_{1})\eta$ and the two black vertical lines mark the end of warmup and stable phases.
While progressive sharpening has been consistently documented at small scales, especially in full-batch settings, its manifestation and relevance at large scales, particularly in the context of online LLM training, remain largely unexplored. Having established that critical sharpness serves as an efficient and reliable proxy for progressive sharpening and Edge of Stability phenomena, we are now in the position to investigate whether progressive sharpening persists in large-scale LLM training. To this end, we analyze the publicly available OLMo-2 77B checkpoints (Walsh et al., 2025). OLMo-2 provides checkpoints throughout both pre-training and mid-training stages, enabling us to study sharpness dynamics at scale without the computational cost of training such models from scratch.
The OLMo-2 models are trained using a two-stage curriculum approach. During the pre-training stage ($90\%$ of compute), the model is trained for 44T tokens on the DCLM dataset (Li et al., 2025), a diverse web-sourced corpus. Following pre-training, the model undergoes a mid-training stage, where it is trained for 5050B tokens on a curated mix. This mix includes academic papers, Wikipedia articles, instruction-following data, StackExchange documents, and approximately $50\%$ of the original pre-training dataset (DCLM). During the pre-training stage, the model is trained using a learning rate schedule consisting of 2,0002{,}000 steps of warmup, followed by a cosine decay down to one-tenth of its peak value [^2] For the OLMo-2 77B models, the cosine decay schedule is designed to reach one-tenth of the peak learning rate at 55T tokens, but is truncated at 44T tokens.. By comparison, during the mid-training stage, the learning rate is linearly decayed to zero. As the learning rate is continuously decreased throughout training after the warmup stage, the EoS threshold increases, and consequently we expect sharpness to progressively increase for the remainder of training. To confirm this, we examine how the critical sharpness evolves across pre-training and mid-training using the publicly available checkpoints (see Section˜8 for the details). Figure˜5 shows that critical sharpness rapidly decreases during early training, but then continually increases (progressive sharpening) throughout both the pre-training and mid-training stages. This result provides the first empirical evidence of progressive sharpening at scale in practical LLM training settings.
Figure 5: Critical Sharpness of OLMo-2 77B exhibits progressive sharpening throughout pre-training and mid-training. In the mid-training figure, the band around the mean trend shows the deviation across the three runs.
4 How much Pre-training data is needed to avoid Catastrophic forgetting?
To go beyond demonstrating pre-existing sharpness phenomena at scale, we next provide a new practical application of critical sharpness, focused on data mixing. We are motivated by the fact that during finetuning, neural networks are prone to catastrophic forgetting, where the model performance degrades on the pretraining dataset and benchmarks as the model adapts to the new task (McCloskey and Cohen, 1989; Luo et al., 2025; McLeish et al., 2025). To mitigate this, several strategies have been proposed (De Lange et al., 2022), with mixing samples from the pre-training data being the most effective (Robins, 1995). This practice is reflected in large-scale training: for example, Walsh et al. (2025) use a mid-training mix consisting of approximately $50\%$ pre-training data. Intuitively, we want to use as much as possible new data in fine-tuning, using as little pre-training rehearsal data as possible to retain base capabilities. However, it remains unclear what fraction of pre-training data is sufficient to effectively prevent catastrophic forgetting. To this end, we leverage critical sharpness to systematically examine the effect of adding pre-training data to the training mix during OLMo mid-training and provide actionable guidance for selecting the pre-training fraction without exhaustive grid search.
The goal of mid-training or finetuning is to improve the performance on specialized domains or tasks while preserving generic capabilities acquired during pre-training. Intuitively, this requires the model to adapt to the new task, while staying within the “pre-training basin” —a region of the parameter space where the pre-training loss remains low. Leaving this basin would be marked by an increase in the pre-training loss, and the critical learning rate quantifies exactly how far we can step before this occurs. To formalize this intuition, we define relative critical learning rate and sharpness, as follows:
Definition 4.1 (Relative Critical learning rate $\eta_{c}^{1\to 2}$ and Sharpness $\lambda_{c}^{1\to 2}$).
Consider a model with parameters $\bm{\theta}$, two loss functions $L_{1}(\bm{\theta})$ and $L_{2}(\bm{\theta})$, and an update direction $\Delta\bm{\theta}$ derived from $L_{2}$. The relative critical learning rate is the smallest learning rate for which taking a step in the direction $\Delta\bm{\theta}_{2}$ increases the loss $L_{1}(\bm{\theta})$:
$\eta_{c}^{1\to 2}=\min_{\eta>0}\{\eta\mid L_{1}(\bm{\theta}-\eta\Delta\bm{\theta}_{2})>L_{1}(\bm{\theta})\}.$ (6)
The corresponding relative critical sharpness is $\lambda_{c}^{1\to 2}=2/\eta_{c}^{1\to 2}$.
This definition is general and applies whenever optimization on one objective may affect performance on another. The two losses can correspond to different tasks (e.g., pre-training vs. finetuning), different loss functions (e.g., next-token prediction vs. reinforcement learning objectives), or different data characteristics (e.g., short vs. long context). In this work, we focus on the finetuning setting, where $L_{1}({\bm{\theta}})$ is the pre-training loss on a general text corpus and $L_{2}({\bm{\theta}})$ is the finetuning loss on a specialized dataset such as math.
To assess the impact of pre-training data fraction in the training mix during fine-training, we consider the OLMo-2 77B pre-trained checkpoint as our starting point. We compute the update direction from a mixture composed of DCLM (pre-training data) and the math subset of Dolmino mix (Walsh et al., 2025), and examine how the relative critical sharpness varies with the DCLM fraction in this mixture [^3] When measuring the relative critical sharpness, we do not update the model parameters..
Figure˜6(a) shows how the relative critical sharpness varies with the DCLM ratio in the training mix for several evaluation tasks. When the mix contains mostly math data (low DCLM ratio), the sharpness for DCLM is an order of magnitude higher than for other tasks, indicating that the pre-training loss landscape is particularly sharp and sensitive to finetuning updates in this regime. As more pre-training data is added, the relative critical sharpness for most tasks decreases, suggesting that the landscapes align. On the other hand, when the DCLM ratio approaches one, the sharpness for downstream tasks such as Math, GSM8K, and MMLU increases, meaning that these tasks become the limiting factor for the maximum stable learning rate. Notably, there is an intermediate DCLM ratio (around 0.7) where the sharpness curves for different tasks intersect, representing a sweet spot that allows for the largest possible learning rate without being constrained by any single task.
Figure 6: (a) Relative critical sharpness (Definition˜4.1) for various evaluation tasks. The shaded region around the mean trends denotes the variation across batches. (b, c) GSM8K and MMLU accuracy as a function of pre-training (DCLM) mix ratio and learning rate. Red indicates a decrease in performance relative to the checkpoint, white indicates no change, and blue indicates an improvement. The black dashed line denotes the smallest critical learning rate among at the pre-trained checkpoint across tasks from (a) and white dashed line denotes the critical learning rate for the corresponding task.
In Figure˜6(b, c), we evaluate the impact of the pre-training mix by training the pre-trained checkpoint for 11B tokens and measuring downstream accuracy on GSM8K and MMLU benchmarks. In these plots, red indicates a decrease in performance relative to the checkpoint, white indicates no change, and blue indicates an improvement. We observe a natural trade-off: GSM8K accuracy is typically maximized outside the pre-training basin (i.e., with low DCLM ratio and high learning rate), but this comes at the expense of MMLU performance, which is best preserved within the pre-training basin. Therefore, if the sole objective is to improve on the finetuning task (math), then training primarily on it with a large learning rate is effective, though it may lead to forgetting on other tasks (MMLU). In contrast, if the goal is to improve on the finetuning task (math) while maintaining pre-training performance (MMLU), it is essential to include pre-training data in the mix, with a sweet spot emerging around a DCLM ratio of $0.6$ and a learning rate of 3e3e-$05$ in our experiment. Remarkably, this is close to the optimum of $0.7$ suggested by the critical sharpness analysis. This result highlights the importance of balancing task-specific and general data in the training mix. Once the pre-trained loss $L_{1}$ increases ($\eta>2/\lambda_{c}^{1\to 2}$), there is no guarantee that further updates will improve $L_{1}$ again, unless the pre-train data is injected into the mix.
In Section˜9.4, we extend the relative critical sharpness analysis to the datamix used in OLMo-2 mid-training. The results are consistent: we observe a sweet spot at a DCLM ratio of $0.6$, close to the $0.5$ ratio used in the original OLMo-2 training. We leave the validation of this prediction through downstream evaluation to future work.
5 Related Works
In neural network training, if the learning rate exceeds the threshold $\sim 2/\lambda^{H}_{\max}$, it causes the loss to increase in the next training step (Wu et al., 2018). However, Kalra and Barkeshli (2023) demonstrated that the empirical critical learning rate can be much higher than this theoretical threshold derived from convex analysis, with critical learning rate reaching up to $40/\lambda^{H}_{\max}$ when sharpness decreases significantly during training. Our work extends theirs by showing that a similarly large critical learning rate can also arise when sharpness increases during training, due to contributions from other eigendirections (Figure˜3). Beyond its role in characterizing training instabilities, Hessian sharpness also exhibits progressive sharpening during training. In particular, Cohen et al. (2021) showed that sharpness increases when the learning rate is decayed, which we corroborate with additional analysis at large scales (Figures˜4 and 5).
The works most closely related to ours are those by Kalra and Barkeshli (2024) and Roulet et al. (2024). Kalra and Barkeshli (2024) used critical learning rate to set the initial learning rate during warmup. By comparison, Roulet et al. (2024) proposed setting the learning rate at ‘edge’ throughout training, i.e., $\eta_{t}=2/\lambda_{\text{dir}}$, and show that it can outperform constant learning rate training, but not typical learning rate schedules, consisting of learning rate warmup followed by decay. Vaswani et al. (2019) also study setting the learning rate as $\eta=1/\lambda_{\text{dir}}$. In contrast, we use the critical learning rate to study the curvature dynamics of large-scale models and examine the effect of pre-training data on catastrophic forgetting.
Catastrophic forgetting is a long-standing problem in neural networks, first identified by McCloskey and Cohen (1989), where model performance degrades on previously learned tasks as it adapts to the new task. Approaches to mitigate catastrophic forgetting can be categorized into (i) regularization methods (Kirkpatrick et al., 2017; Ahn et al., 2019), (ii) ensembling and parameter isolation (Rusu et al., 2016; Aljundi et al., 2017), and (iii) rehearsal (Robins, 1995; Lopez-Paz and Ranzato, 2017). Among these, rehearsal —mixing samples from the pre-training data —has become the most widely adopted strategy due to its simplicity and effectiveness. Recent works have demonstrated that catastrophic forgetting persists in LLM finetuning (Luo et al., 2025; Huang et al., 2024; Scialom et al., 2022). Particularly related to our work is that of Chen et al. (2025), who show that if finetuning retains previously learned capabilities, the model remains in “most-case” and “worst-case” basins.
6 Discussion and Conclusion
In this work, we analyzed critical sharpness, a computationally efficient measure for studying the training dynamics of LLMs. Our results demonstrate that critical sharpness reliably captures key Hessian sharpness phenomena such as progressive sharpening and Edge of Stability, requiring fewer than $10$ forward passes, avoiding challenges associated with Hessian-based methods. Using this measure, we provided the first empirical evidence of progressive sharpening at the 77B parameter scale.
We also introduced relative critical sharpness, which quantifies the curvature of one loss landscape along the update direction of another. Using this measure, we identified a sweet spot in the pre-training data fraction that balances specialization and retention during finetuning, enabling practitioners to evaluate data composition choices without extensive ablations. Beyond finetuning, relative critical sharpness provides a general framework for analyzing changes in the loss landscape due to distribution shifts, changes in loss functions, or modifications to the training data mixture.
We believe critical sharpness can extend beyond the settings studied here. More broadly, our results demonstrate that scalable curvature measures can provide actionable insights for large-scale training, from understanding optimization dynamics to informing data composition decisions.
Acknowledgements
We would like to thank Tianyu He and Darshil Doshi for helpful discussions and detailed comments on the manuscript, and Sean McLeish, Konstantin Mishchenko, Aaron Defazio, Maissam Barkeshli, Benjamin Therien, and Jesse Dodge for helpful discussions.